Fermi National Laboratory


Accelerator Update

Wednesday July 14

The day shift began with Operations monitoring store 3653 and the Recycler stash, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

At 8:35 AM, Operations reported noticing that shortly after the Tevatron (TeV) RF station (TRF8) tripped off at 6 AM this morning, the Tevatron Electron Lens (TEL) began putting out less current. An expert investigated.

Operations turned off the Booster beam at 9:04 AM to bring the power supply for BRF8 back on line. Beam resumed at 9:27 AM.

A Main Injector (MI) expert began a slipstacking study at 9:31 AM.

At 11:52 AM, MI RF stations 1, 7, & 11 tripped off. They all reset okay.

Operations reported that there was a vacuum burst in the A150 transfer line a little while ago. They reported at 12:01 PM that everything was fine.

A Booster expert began a damper study at 5:08 PM.

Operations terminated store 3653 at 8:01 PM.

TeV RF expert began work on TRF8 and a TEL expert investigated problems at 8:07 PM. Neither expert could find anything wrong with their devices. Both were back up and running by 9:37 PM.

At 10:16 PM, Operations reported problems with the TEL. Another expert came in to help.

Operations reported at 10:45 PM that they were starting the TeV shot setup while the TEL experts continued their work.

The TEL experts reported at 11:18 PM that they had replaced a power supply and that the TEL was up and running.

Thursday July 15

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash, and with MiniBooNE taking beam.

Operations established store 3655 at 12:46 AM with an initial luminosity of 77.97E30.

Operations began stashing antiprotons into the Recycler at 3:31 AM. They completed the shots by 4:38 AM.

At 7:45 AM, two Booster RF stations (BRF7 & 8) tripped off due to high vacuum pressure. They both reset okay.

A Pbar expert began a study period at 11:13 AM.

Operations established beam to the Switchyard (SY) dump at 2:02 PM. Experts then tuned beam to the Meson.

A TeV expert raised separator voltages at 2:05 PM.

A MI expert began a ramp study at 2:23 PM. She was done an hour and fifteen minutes later.

At 4:50 PM, Operations reported a problem with SY power supplies V210 and Q211. Operators called in power supply experts.

A MI expert began another ramp study at 8:58 PM.

An MTest experimenter reported that they were through with beam for the night. Operations disabled their beamline.

At 9:56 PM, the D-Zero (D0) experiment called to complain about the colliding beam. While an operator was talking with a TeV expert, the problem got better. The expert suggested adjusting the tunes and chromaticities.

The power supply technicians had SY V210 running by 11:24 PM. However, Q211 hadn't been commissioned yet, and needed official permission to run. Operations established beam to MCenter.

Friday July 16

The midnight shift began with Operations monitoring store 3655 and the stash, with Pbar stacking, and with the MiniBooNE and MCenter experiments taking beam.

TeV experts began an end-of-store study shortly after 5:37 AM.

Operations terminated store 3655 at 6:49 AM.

Operations prepared the Collider Detector at Fermilab (CDF) and D0 collision halls for access at 7 AM.

Plans for Friday and this Weekend

The plan for this morning is to establish a Mixed Antiproton Mode store. The plans for the weekend are to continue with stacking and storing.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 7/16/2004   email Fermilab